# The XMAGNET exascale MHD simulations of SMBH feedback in galaxy groups and clusters: Overview and preliminary cluster results

### v1

@@ 4.1 Cluster self-regulation
[…]
Conversely, some studies have shown that jet precession or reorientation might play a role in the dynamics of the cold phase. In Beckmann et al. (2019), rapid re-orientation of the jets on timescales of a few tens of Myr pushes the cold phase outwards in a nearly isotropic way, effectively shutting off the fueling of the AGN for a few tens of Myr and possibly resulting in the alternation between heating and cooling dominated phases.
@@ 4.6.3 Stellar Feedback
The algorithms that we use to approximate star formation and supernova mass and energy feedback in the central cluster galaxy (Sec. 2.2.11) are relatively simple, and in particular are volumetric rather than particle-based in nature. In addition, feedback that approximates Type II supernovae originating from cold, dense gas in the halo center is approximated as instantaneous feedback rather than including a physically-appropriate delay. We include feedback using this volumetric method because both theory and previous simulations have shown that it is potentially important for the self-regulation of the circumgalactic media in massive galaxies (see, e.g., Voit et al., 2020; Prasad et al., 2020). Still stellar feedback is globally sub-dominant. It is unclear if having a more granular stellar feedback algorithm would make a significant difference at the spatial and mass resolutions we are considering (in contrast to, e.g., at the dwarf or Milky Way-galaxy scale). This is also an area of potential future study.
@@ 4.7 Future papers in this series
This paper is the first in a series of publications presenting the results from our simulation campaign, and presents a high-level overview of a subset of the simulations. In the following papers we will perform detailed analyses of specific aspects of the simulations and examine a broader range of galaxy groups and clusters than is presented in this paper. Specifically, we plan to do the following:
@@ 5 Conclusions and Outlook
The authors thank Greg Bryan, Hui Li, Yuan Li, Ricarda Beckmann, and Rainer Weinberger for insightful discussions and the Oak Ridge Leadership Computing Facility support staff, in particular John Holmen, for incredibly valuable support at all stages of the INCITE project that resulted in this and following publications. BDW thanks Michael Zingale for useful discussions of performance, scaling, and bugs on Frontier. BWO acknowledges support from NSF grants #1908109 and #2106575, NASA ATP grants NNX15AP39G and 80NSSC18K1105, and NASA TCAN grant 80NSSC21K1053. BWO also thanks the Los Alamos National Laboratory Center for Nonlinear Studies for support via the Ulam Sabbatical Fellowship.
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. These resources were provided by as part of the DOE INCITE Leadership Computing Program under allocation AST-146 (PI: Brian O’Shea).
MF and MB acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. DP is supported by the Royal Society through UKRI grant RF-ERE-210263 (PI: Freeke van de Voort). All simulations were performed using the public MHD code AthenaPK, which makes use of the Kokkos library (Trott et al., 2022) and the Parthenon adaptive mesh refinement framework (Grete et al., 2023). All data analysis was performed with yt (Turk et al., 2011, 2024), Matplotlib (Hunter, 2007), and Numpy (Harris et al., 2020). We thank their authors for making these software and packages publicly available.

### v2

@@ 4.1 Cluster self-regulation
Our results differ from previous results of self-regulating galaxy clusters. Hydrodynamic simulations of idealized clusters including cold gas-triggered AGN, such as those performed by Gaspari et al. (2012); Prasad et al. (2015); Li et al. (2015) and Meece et al. (2017), typically show a greater variability in AGN energies and large-scale cyclic behavior on $\sim$ Gyr time scales. On the other hand, Ehlert et al. (2023) perform simulations of self-regulated magnetized AGN feedback in cool-core clusters and see behavior that is more similar to our results – i.e., the jets in their simulations, regardless of the precise details of jet momentum density or triggering mechanism, are continuously on. An interesting intermediate case is provided by Yang & Reynolds (2016), who use cold-gas triggered hydrodynamical jets in an initially magnetized medium. This work shows highly variable AGN energies but no indication of Gyr-scale patterns. It is unclear whether the difference in jet behavior between hydrodynamical and magnetohydrodynamical simulations of self-regulated AGN feedback is primarily caused by the magnetic fields themselves or by other feedback parameters. For example, the momentum density of the jet affects the jet behavior – see, e.g., the discussion in Section 7.2 of Prasad et al. (2020), Section 5.1 of Prasad et al. (2022), and Section 3 of Weinberger et al. (2023). In general, lighter jets are more easily contained at smaller cluster-centric radii and are also more easily deflected by cold gas than heavier jets. This may result in the simulated cluster’s center displaying less variability in its thermodynamic state than for heavier jets.
Conversely, some studies have shown that jet precession or reorientation might play a role in the dynamics of the cold phase. In Beckmann et al. (2019), rapid reorientation of the jets on timescales of a few tens of Myr pushes the cold phase outwards in a nearly isotropic way, effectively shutting off the fueling of the AGN for a few tens of Myr and possibly resulting in the alternation between heating and cooling dominated phases.
@@ 4.6.3 Stellar Feedback
The algorithms that we use to approximate star formation and supernova mass and energy feedback in the central cluster galaxy (Sec. 2.2.11) are relatively simple, and in particular are volumetric rather than particle-based in nature. In addition, feedback that approximates Type II supernovae originating from cold, dense gas in the halo center is approximated as instantaneous feedback rather than including a physically-appropriate delay. We include feedback using this volumetric method because both theory and previous simulations have shown that it is potentially important for the self-regulation of the circumgalactic media in massive galaxies (see, e.g., Voit et al., 2020; Prasad et al., 2020). Still stellar feedback is globally sub-dominant. It is unclear if having a more granular stellar feedback algorithm, which might impact the cold gas dynamics in the central region and, thus, the AGN fueling, would make a significant difference at the spatial and mass resolutions we are considering (in contrast to, e.g., at the dwarf or Milky Way-galaxy scale). This is also an area of potential future study.
@@ 4.7 Future papers in this series
This paper is the first in a series of publications presenting the results from our simulation campaign, and presents a high-level overview of a subset of the simulations. In the following papers we will perform detailed analyses of specific aspects of the simulations and examine a broader range of galaxy groups and clusters than is presented in this paper. First results on the velocity structure functions (VSFs) of the multiphase ICM and the effects of projections (at various viewing angles and times) have already been published in Fournier et al. (2025). Key findings include that there is no correlation between the hot and cold phase VSFs and that that observational biases such as projection effects, atmospheric seeing and the viewing angle cannot be ignored when interpreting line-of-sight VSF. In addition, we plan to do the following:
@@ 5 Conclusions and Outlook
The authors thank Greg Bryan, Hui Li, Yuan Li, Ricarda Beckmann, and Rainer Weinberger for insightful discussions and the Oak Ridge Leadership Computing Facility support staff, in particular John Holmen, for incredibly valuable support at all stages of the INCITE project that resulted in this and following publications. BDW thanks Michael Zingale for useful discussions of performance, scaling, and bugs on Frontier. PG acknowledges the support by the DFG grant SCHW 1358/5-1 within the priority program SPP 1992. BWO acknowledges support from NSF grants #1908109 and #2106575, NASA ATP grants NNX15AP39G and 80NSSC18K1105, and NASA TCAN grant 80NSSC21K1053. BWO also thanks the Los Alamos National Laboratory Center for Nonlinear Studies for support via the Ulam Sabbatical Fellowship. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. These resources were provided by as part of the DOE INCITE Leadership Computing Program under allocation AST-146 (PI: Brian O’Shea).
MF and MB acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). DP is supported by the Royal Society through UKRI grant RF-ERE-210263 (PI: Freeke van de Voort). All simulations were performed using the public MHD code AthenaPK, which makes use of the Kokkos library (Trott et al., 2022) and the Parthenon adaptive mesh refinement framework (Grete et al., 2023). All data analysis was performed with yt (Turk et al., 2011, 2024), Matplotlib (Hunter, 2007), and Numpy (Harris et al., 2020). We thank their authors for making these software and packages publicly available.